Technical Observing AGW1 & AGW2 Testing 20131214 UT

Observers: DMiller (Tucson), DThompson (LBTO)
Telescope Operator: GBechetti (LBTO)
AO Support: DMiller (Tucson)
Software Support: MDeLaPena (Tucson)
Telescope Support: MMidkiff (LBTO)
Instrument Support: ESolheid (LBTO)
Engineering Support: DAshby (Tucson)

Summary

Details

11:15 Start by measuring on-axis probe position for AGw1 and AGw2

11:20 use setxy to put probes at 0,612.5

11:21 use readGuideCam left/right -e 90000
  • images in /home/dlmiller/data/20131214/center_(lfbg/rfbg)
  • image 1

11:30 LEFT
  • center at 120, 95.4
  • full frame center at 255,231
  • guider hotspot at 255, 230
  • Perfect no change

11:30 RIGHT
  • center at 126.5, 95
  • full frame center at 270, 205
  • guider hotspot at 269, 265 (note, checked that SH spots were in center of subaps. Hotspot is correct)
  • change /lbt/oacontrol/etc/oacontrol.conf
    • # JMH/DLM 2013-10-25
    • #ooffset=419522,&1Q37
    • # move probe up 60 pixels DLM 2013-12-14
    • ooffset=417722,&1Q37

11:35 problems with cable chain. Geno had to reset on telescope

12:10 Send preset. AGw2 communication problem. Geno home stages. OK

12:19 Geno stop and start oacserver.

12:21 send preset OK

12:30 use readGuideCam left/right -e 90000
  • images in /home/dlmiller/data/20131214/center_(lfbg/rfbg)
  • image 2
  • FORGOT to move guide probes to 0,0

12:35 use readGuideCam left/right -e 90000
  • images in /home/dlmiller/data/20131214/center_(lfbg/rfbg)
  • image 3

12:40 LEFT
  • center at 120, 97
  • Perfect no change

12:40 RIGHT
  • center at 124, 94
  • no change. config change not applied

12:49 RIGHT
  • make ooffset of -1800 microns to oacontrol.conf.default
  • Geno restart oacserver

12:51 preset

13:00 AGw2 now is almost correct position
  • image 5

13:05 trouble with rotator and the preset got STUCK. Had to stop/start GCS to get preset to cancel

13:25 put oacontrol.conf.default back to original value. Geno stop/start oacserver. See if .default file is really in control of AGw2.

13:29 preset

13:30 Sky too bright. Close dome

13:43 Set oacontrol.conf.default back to working value. Instead of -1800 I used -1700 because hotspot was about 5 pixels too high. Will have to test on-sky again.

-- DougMiller - 14 Dec 2013
Topic revision: r2 - 14 Dec 2013, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback